OpenStudy (anonymous):

A grocer bought a box of soap containing 72 bars for $24.50. Some of the soap he sold at 3 bars for $1.25, and the rest at $.35 a bar. This gave a profit of $3.90. How many bars did he sell at each price?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

The total amount he spent on purchasing 72 bars = 24.50 Lets say he sold x bars at 1.25 and y bars at .35. So the total number of bars: \[x+y=24\] The total amount he made by selling the bars: \[1.25\times x + .35\times y = 24.50+3.90 = 28.40\] Solve the two equations for x and y and simplify.

OpenStudy (anonymous):

magarwal you second expression is correct. but te first expression does not make sense. it should be 3x + y = 72 where x is the number of batches sold at 1.25 and y is the number of bars sold individually

OpenStudy (anonymous):

when you solve simultaneously you get x=16 and y=24. i.e he sold 16 batches at 1.25 and 24 individuals at .35 each

OpenStudy (anonymous):

I am sorry, my bad the first expression should have been as KBOY said \[3x+y=72\]